18) You own 10 futures contracts on gold that you purchased at a quoted price of 1,483.9. The current price quote is 1,403.2. The contract size is 100 ounces and the quotes are expressed in dollars and cents per ounce. What is your current profit or loss on this investment?

19) Alicia owns 600 shares of Danube stock. She thinks the market price will continue to rise but would like to ensure that she can get at least $47.50 a share should she decide to sell her shares. The 47.50 call option is quoted at $1.05 bid, $1.11 ask. The 47.50 put is quoted at $0.80 bid, $0.89 ask. How much will it cost her to ensure that she can sell all of her shares for at least $47.50 each?

20) The Stable Utility Fund has an offering price of $54.35 and an NAV of $52.19. What is the front-end load percentage?
Answer #1

18) Answer: Loss $80,700

10 future contracts

contract size = 100 ounce

Purchase price = 1483.9
Current Price = 1403.2

Purchase value = 1483.9 * 10 * 100 =14,83,900

Current Value = 1403.2 * 10 * 100 = 14,03,200

Current loss = $80,700

19) Answer = $534

owns 600 shares

To ensure that she can sell all of her shares for at least $47.50 each, she need to purchase 47.50 put option at price of $0.89.

Cost of buying put option = 600 * $0.89 = $534

To ensure that she can sell all of her shares for at least $47.50, the cost will be $534.

20) Answer = 3.9742%

offer price = NAV + front-end load

Front-end load = offer price - NAV

Front-end load = $54.35 - $52.19 = $2.16

Front-end load percentage = 1 - (NAV/offer price)

Front-end load percentage = 1-(52.19/54.35) = 0.039742 = 3.9742%